# Open Packing in Graphs: Bounds and Complexity

### v1

@@ Abstract
It is known that Open Packing is NP-complete on split graphs (i.e., the class of $\{2K_{2},C_{4},C_{5}\}$-free graphs) [Ramos et al., 2014]. In this work, we complete the study on the complexity (P vs NPC) of Open Packing on $H$-free graphs for every graph $H$ with at least three vertices by proving that Open Packing is (i) NP-complete on $K_{1,3}$-free graphs and (ii) polynomial time solvable on $(P_{4}\cup rK_{1})$-free graphs for every $r\geq 1$. In the course of proving (ii), we show that for every $t\in\{2,3,4\}$ and $r\geq 1$, if $G$ is a $(P_{t}\cup rK_{1})$-free graph, then $\rho^{o}(G)$ is bounded above by a linear function of $r$. In addition, we find near-optimal upper bounds for the total domination number in the class of $(P_{t}\cup rK_{1})$-free graphs for every $t\in\{3,4\}$ and $r\geq 1$. Moreover, we show that Open Packing parameterized by solution size is W[1]-complete on $K_{1,3}$-free graphs and Max-Open Packing is hard to approximate within a factor of $(n^{\frac{1}{2}-\delta})$ for any $\delta>0$ on $K_{1,3}$-free graphs unless $P=NP$. Further, we prove that Open Packing is (a) NP-complete on $K_{1,4}$-free split graphs and (b) polynomial time solvable on $K_{1,3}$-free split graphs. We prove a similar dichotomy result on split graphs with degree restrictions on the vertices in the independent set of the clique-independent set partition of the split graphs.
@@ 1 Introduction
In a graph $G(V,E)$, a vertex subset $D$ of $G$ is called a total dominating set in $G$ if every vertex in $V(G)$ is adjacent to some vertex in $D$. In other words, $V(G)=\cup_{u\in D}N_{G}(u)$, where $N_{G}(u)$ denotes the set of vertices in $V(G)$ that are adjacent to $u$ in $G$. This implies that if $D$ is a total dominating set in $G$, then $|D\cap N_{G}(u)|\geq 1$ for every vertex $u\in V(G)$. Note that by the definition of a total dominating set, it is evident that a graph $G$ admits a total dominating set if and only if $G$ has no isolated vertices. The cardinality of a smallest total dominating set in $G$ is called the total domination number, $\gamma_{t}(G)$, of $G$. The total domination problem is useful in facility location problems, monitoring computer networks and electronic communications [7]. In this article, we study the total dominating set problem and its dual problem called the open packing in graphs [8]. A vertex subset $S$ of a graph $G$ is called an open packing in $G$ if for every pair of distinct vertices $x,y\in S$, $N_{G}(x)\cap N_{G}(y)$ is empty. The open packing number of $G$ denoted by $\rho^{o}(G)$ is the size of a largest open packing in $G$. Note that if $S$ is an open packing in $G$, then $|S\cap N_{G}(u)|\leq 1$ for every $u\in V(G)$. This validates the primal-dual relation between the total dominating set and the open packing, and thus $\rho^{o}(G)\leq\gamma_{t}(G)$. The following are the decision and the optimisation versions of the problems total dominating set and open packing.
@@
Total dominating set is one of the well-studied problems in the literature and an extensive list of results can be found in [5, 7, 10]. In the context of our research, it would be interesting to note that Total Dominating Set is NP-complete on $K_{1,5}$-free split graphs [24] and polynomial time solvable on $K_{1,4}$-free split graphs [20]. Also, Total Dominating Set is NP-complete on $K_{1,3}$-free graphs [14] and an optimal total dominating set of a chordal bipartite graph can be found in polynomial time [3]. The study on open packing of graphs was initiated by Henning and Slater [9] and Rall [18] proved that for every non-trivial tree $T$, $\gamma_{t}(G)=\rho^{o}(G)$. In a recent work [21] (yet to be published), we extended this result by proving that the total domination number and the open packing number are equal when the underlying graph is a chordal bipartite graph with no isolated vertices. It is also known that Open Packing is NP-complete (NPC in short) on split graphs (equivalently, the class of $\{2K_{2},C_{4},C_{5}\}$-free graphs) [19] and planar bipartite graphs of maximum degree three (a subclass of $K_{3}$-free graphs) [21]. In this work, we complete the study of Open Packing on $H$-free graphs for every graph $H$ with at least three vertices by proving the following theorems.
@@
We also show that the bounds on $\rho^{o}(G)$ given in (ii) are tight. Moreover, we use the bounds on the total domination number in (a) and (b) to show that the results equivalent to those of Theorems 1 and 2 hold for Total Dominating Set in $H$-free graphs as well (see Theorem 9). This implies that Total Dominating Set and Open Packing are of the same nature in the view of classical complexity (P vs NPC) in $H$-free graphs for every graph $H$. A comparative study of complexity (P vs NPC) between Total Dominating Set and Open Packing in certain classes of graphs is given in Table 1. The rest of the results in this article are motivated by Table 1 and the question: ‘Does there exist a graph class where these two problems ‘differ’ from each other in the view of classical complexity (P vs NPC)?’
@@
We answered this question in affirmative (see Table 2) with the following list of dichotomy results in subclasses of split graphs.
@@
    Open Packing is NP-complete on $K_{1,r}$-free split graphs for $r\geq 4$ and is polynomial time solvable for $r\leq 3$.
@@ 2 Preliminaries
We follow West [23] for terminology and notation. The graphs considered in this work are simple and undirected unless specified otherwise. Given a graph $G(V,E)$, let $n$ and $m$ denote the number of vertices and the number of edges in $G$, respectively. Given a vertex $x\in V(G)$, the (open) neighbourhood of $x$ in $G$ is defined as $N_{G}(x)=\{y\in V(G)\,:\,xy\in E(G)\}$ and let the degree of a vertex $x$ in $G$ be $deg_{G}(x)=|N_{G}(x)|$. The closed neighbourhood of a vertex $x$ in $G$ is defined as $N_{G}[x]=\{x\}\cup N_{G}(x)$. We use $N(x)$, $N[x]$ and $deg(x)$ instead of $N_{G}(x)$, $N_{G}[x]$ and $deg_{G}(v)$ respectively, when there is no ambiguity on $G$. A vertex $x$ in $G$ is called an isolated vertex in $G$, if $N_{G}(x)=\emptyset$. Given $U\subseteq V(G)$, the subgraph of $G$ induced by $U$ is denoted as $G[U]$. Given a graph $H$, $G$ is said to be $H$-free if no induced subgraph of $G$ is isomorphic to $H$. For a vertex $x\in V(G)$, let $E_{G}(x)$ denote the set of all edges incident on $x$, and for an edge $e\in E(G)$, let $V_{G}(e)$ denote the end vertices of $e$ in $G$. Note that for $u\in V(G)$ and $e\in E(G)$, the edge $e\in E_{G}(u)$ if and only if $u\in V_{G}(e)$. For a graph $G$, the line graph $L(G)$ of $G$ is a graph with $V(L(G))=E(G)$ and two entries $e,e^{\prime}\in V(L(G))$ are adjacent in $L(G)$ if and only if $V_{G}(e)\cap V_{G}(e^{\prime})\neq\emptyset$. Given two graphs $H$ and $H^{\prime}$, the graph union $H\cup H^{\prime}$ is defined as $V(H\cup H^{\prime})=V(H)\cup V(H^{\prime})$ and $E(H\cup H^{\prime})=E(H)\cup E(H^{\prime})$. If $V(H)$ and $V(H^{\prime})$ are not disjoint, then we call the graph union $H\cup H^{\prime}$ as merger of the graphs $H$ and $H^{\prime}$. For $p\in\mathbb{N}\cup\{0\}$ and a graph $H$, the graph $pH$ is defined as the union of $p$ disjoint copies of $H$. Given two disjoint graphs $G$ and $H$, we say that (i) a vertex $x$ of the graph $G$ is replaced by the graph $H$ if a graph $G^{\prime}$ is constructed with $V(G^{\prime})=(V(G)\setminus\{x\})\cup V(H)$ and $E(G^{\prime})=(E(G)\setminus\{xy:y\in V(G)\text{ and }xy\in E(G)\})\cup E(H)\cup\{wy:w\in V(H),y\in(V(G)\setminus\{x\})\text{ and }xy\in E(G)\}$, and (ii) a vertex $x$ of a graph is identified with a vertex $y$ of the graph $H$, if the vertex $y$ in $H$ is relabelled as $x$ and the graphs $G$ and $H$ are merged.
@@
A vertex subset $C$ of $G$ is called a clique in $G$ if every pair of distinct vertices in $C$ is adjacent in $G$. A vertex subset $I$ of $G$ is called an independent set in $G$ if no pair of vertices in $I$ are adjacent in $G$. The size of a largest independent set in $G$ is called the independence number of $G$, and is denoted by $\alpha(G)$. Let $P_{n}$, $C_{n}$ and $K_{n}$ denote the path, cycle and complete graph on $n$ vertices, respectively. A set $D\subseteq V(G)$ is called a dominating set in $G$, if every vertex in $V(G)\setminus D$ is adjacent to some vertex in $D$. An edge subset $F$ of a graph $G$ is called a matching in $G$, if $V_{G}(e)\cap V_{G}(e^{\prime})=\emptyset$ for every pair of distinct $e,e^{\prime}\in F$.
@@ 3 $H$-free Graphs
We dedicate this section to prove the dichotomy result on the open packing problem stated in Theorem 2. Observation 1 helps us to prove the necessary part of Theorem 2.
@@
For $p\geq 4$, let $H$ be a graph on $p$ vertices such that $H\notin\{P_{4}\cup(p-4)K_{1},P_{3}\cup(p-3)K_{1},K_{2}\cup(p-2)K_{1},pK_{1}\}$. Then, $H$ contains at least one of $K_{3},2K_{2},C_{4},K_{1,3}$ or $C_{5}$ as an induced subgraph.
@@
Open Packing is NP-complete on (i) $2K_{2}$-free graphs (ii) $C_{4}$-free graphs, (iii) $C_{5}$-free graphs and (iv) triangle-free graphs because of the facts that Open Packing is NP-complete on (a) split graphs (i.e., the class of $\{2K_{2},C_{4},C_{5}\}$-free graphs) [19] and (b) bipartite graphs (a subclass of $K_{3}$-free graphs) [21, 22] and (c) Open Packing is in NP in every subclass of simple graphs.
Next, we prove the following results on $K_{1,3}$-free graphs: (i) Open Packing is NP-complete, (ii) Max-Open Packing is hard to approximate within a factor of $(N^{(\frac{1}{2}-\delta)})$ for any $\delta>0$ unless P = NP (where $N$ denotes the number of vertices in a $K_{1,3}$-free graph) and (iii) Open Packing parameterized by solution size is W[1]-complete.
@@ 3.1 $K_{1,3}$-free Graphs
Step 1 : Replace each edge $e=uu^{\prime}$ in $G$ by a three vertex path $ueu^{\prime}$ in $G^{\prime}$. Step 2 : For every $u\in V(G)$, make $E_{G}(u)$ a clique in $G^{\prime}$. Step 3 : For every vertex $u_{i}\in V(G)$ with exactly one edge, say $e$ incident on it in $G$, introduce a vertex $v_{i}$ and two edges $u_{i}v_{i},v_{i}e$ in $G^{\prime}$.
@@
We can observe that in the graph $G^{\prime}$, for every vertex $\text{for every vertex }u_{i}\in V(G)$
@@
The above equations imply that by Steps 2 and 3 of Construction 1, $N_{G^{\prime}}(u_{i})$ and $N_{G^{\prime}}(v_{i})$ are cliques in $G^{\prime}$. Also, by Construction 1 for every $e\in E(G)$ with $V_{G}(e)=\{u,u^{\prime}\}$,
@@
This implies that $N_{G^{\prime}}(e)$ is a union of two cliques in $G^{\prime}$. We know that a graph $H$ is $K_{1,3}$-free if and only if $N_{H}(x)$ is a union of at most two cliques for all $x\in V(H)$. So, Equations (1), (2) and (3) imply that $G^{\prime}$ is $K_{1,3}$-free.
Next, we prove Claims 1-5 which are helpful to complete the proof of guarantee of Construction 1.
@@
Let $H$ be a graph and let $U=\{x,y,z\}$ be a vertex subset of $H$ that induces a triangle in $H$. Then, for any open packing $S$ in $H$, $|S\cap U|\leq 1$.
@@
Let $u_{i}\in V(G)$ such that $deg_{G}(u_{i})=1$ and let $S$ be an open packing in $G^{\prime}$ with $v_{i}\in S$. Then, (i) $u_{i}\notin S$ and (ii) $S^{\prime}=(S\setminus\{v_{i}\})\cup\{u_{i}\}$ is an open packing in $G^{\prime}$ and (iii) $|S|=|S^{\prime}|$.
Since $deg_{G}(u_{i})=1$, there exists an edge $e\in E(G)$ such that $E_{G}(u_{i})=\{e\}$. Then, $\{u_{i},v_{i},e\}$ induces a triangle in $G^{\prime}$ with $N_{G^{\prime}}(u_{i})=\{e,v_{i}\}$ and $N_{G^{\prime}}(v_{i})=\{e,u_{i}\}$. Therefore by Observation 2, for any open packing $S_{1}$ in $G^{\prime}$, $|S_{1}\cap\{u_{i},v_{i},e\}|\leq 1$. Thus, $v_{i}\in S$ implies that $u_{i},e\notin S$. This proves (i) of Claim 2. Next, we prove (ii). Since $N_{G^{\prime}}(u_{i})=\{e,v_{i}\}$ and $N_{G^{\prime}}(v_{i})=\{e,u_{i}\}$, for any vertex $x\in S\setminus\{v_{i}\}$, $x$ and $v_{i}$ have no common neighbour in $G^{\prime}$ implies that $x$ and $u_{i}$ have no common neighbour in $G^{\prime}$. This together with the fact that $S\setminus\{v_{i}\}$ is an open packing in $G^{\prime}$ implies that $S^{\prime}=(S\setminus\{v_{i}\})\cup\{u_{i}\}$ is an open packing in $G^{\prime}$. Note that (iii) follows from (i).
@@
Let $S$ be an open packing in $G^{\prime}$ with $e=u_{\ell}u_{j}\in S\cap E(G)$ for some $1\leq\ell<j\leq n$. Then, (i) $u_{\ell},u_{j}\notin S$, and (ii) $S_{\ell}=(S\setminus\{e\})\cup\{u_{\ell}\}$ and $S_{j}=(S\setminus\{e\})\cup\{u_{j}\}$ are open packings in $G^{\prime}$ and (iii) $|S|=|S_{\ell}|=|S_{j}|$.
Since $e\in S$, $u_{\ell},u_{j}\notin S$ because of Claim 3. This proves (i). Also, if $v_{\ell}\in V(G^{\prime})$, then $v_{\ell}\notin S$ because of Observation 2 and the fact that $\{v_{\ell},e,u_{\ell}\}$ induces a triangle in $G^{\prime}$. Similarly, if $v_{j}\in V(G^{\prime})$, then $v_{j}\notin S$. Next, we prove that $S_{\ell}=(S\setminus\{e\})\cup\{u_{\ell}\}$ is an open packing in $G^{\prime}$. Note that $S\setminus\{e\}$ is an open packing in $G^{\prime}$. So, if $S_{\ell}$ is not an open packing in $G$, then there exists a vertex $z\in S\setminus\{e\}$ such that $N_{G^{\prime}}(z)\cap N_{G^{\prime}}(u_{\ell})\neq\emptyset$. Let $y\in N_{G^{\prime}}(z)\cap N_{G^{\prime}}(u_{\ell})$. Note that $zu_{\ell}\notin E(G^{\prime})$, else $u_{\ell}\in N_{G^{\prime}}(z)\cap N_{G^{\prime}}(e)$ which would be a contradiction to $e,z\in S$ (an open packing in $G^{\prime}$). Similarly, $zu_{j}\notin E(G^{\prime})$. If $y=e$, then since $e=y\in N_{G^{\prime}}(z)$, we have $z\in(N_{G^{\prime}}(e)\setminus(N_{G^{\prime}}[u_{j}]\cup N_{G^{\prime}}[u_{\ell}]))$, a contradiction to Equation (3) (see Fig. 2). Therefore, $y\neq e$. This implies that $ye\in E(G^{\prime})$, since $y,e\in N_{G^{\prime}}(u_{\ell})$, which is a clique in $G^{\prime}$. Then, $y\in N_{G^{\prime}}(z)\cap N_{G^{\prime}}(e)$, a contradiction to $S$ being an open packing in $G^{\prime}$. Therefore, $S_{\ell}=(S\setminus\{e\})\cup\{u_{\ell}\}$ is an open packing in $G^{\prime}$. A similar argument holds for $S_{j}$. Also, note that (iii) follows from (i).

### v2

@@ Abstract
It is known that Open Packing is NP-complete on split graphs (i.e., the class of $\{2K_{2},C_{4},C_{5}\}$-free graphs) [Ramos et al., 2014]. In this work, we complete the study on the complexity (P vs NPC) of Open Packing on $H$-free graphs for every graph $H$ with at least three vertices by proving that Open Packing is (i) NP-complete on $K_{1,3}$-free graphs and (ii) polynomial time solvable on $(P_{4}\cup rK_{1})$-free graphs for every $r\geq 1$. In the course of proving (ii), we show that for every $t\in\{2,3,4\}$ and $r\geq 1$, if $G$ is a $(P_{t}\cup rK_{1})$-free graph, then $\rho^{o}(G)$ is bounded above by a linear function of $r$. In addition, we find near-optimal upper bounds for the total domination number in the class of $(P_{t}\cup rK_{1})$-free graphs for every $t\in\{3,4\}$ and $r\geq 1$. Moreover, we show that Open Packing parameterized by solution size is W[1]-complete on $K_{1,3}$-free graphs, and Max-Open Packing is hard to approximate within a factor of $(n^{\frac{1}{2}-\delta})$ for any $\delta>0$ on $K_{1,3}$-free graphs unless P = NP. Further, we prove that Open Packing is (a) NP-complete on $K_{1,4}$-free split graphs and (b) polynomial time solvable on $K_{1,3}$-free split graphs. We prove a similar dichotomy result on split graphs with degree restrictions on the vertices in the independent set of the clique-independent set partition of the split graphs.
@@ 1 Introduction
In a graph $G(V,E)$, a vertex subset $D$ of $G$ is called a total dominating set in $G$ if every vertex in $V(G)$ is adjacent to some vertex in $D$. In other words, $V(G)=\cup_{u\in D}N_{G}(u)$, where $N_{G}(u)$ denotes the set of vertices in $V(G)$ that are adjacent to $u$ in $G$. This implies that if $D$ is a total dominating set in $G$, then $|D\cap N_{G}(u)|\geq 1$ for every vertex $u\in V(G)$. Note that by the definition of a total dominating set, it is evident that a graph $G$ admits a total dominating set if and only if $G$ has no isolated vertices. The cardinality of a smallest total dominating set in $G$ is called the total domination number, $\gamma_{t}(G)$, of $G$. The total domination problem is useful in facility location problems, monitoring computer networks, and electronic communications [7]. In this article, we study the total dominating set problem and its dual problem, called open packing in graphs [8]. A vertex subset $S$ of a graph $G$ is called an open packing in $G$ if, for every pair of distinct vertices $x,y\in S$, $N_{G}(x)\cap N_{G}(y)$ is empty. The open packing number of $G$, denoted by $\rho^{o}(G)$, is the size of a largest open packing in $G$. Note that if $S$ is an open packing in $G$, then $|S\cap N_{G}(u)|\leq 1$ for every $u\in V(G)$. This validates the primal-dual relation between the total dominating set and the open packing, and thus $\rho^{o}(G)\leq\gamma_{t}(G)$. The decision and the optimisation versions of the problems of total dominating set and open packing are as follows.
@@
Total dominating set is one of the well-studied problems in the literature and an extensive list of results can be found in [5, 7, 10]. In the context of our research, it would be interesting to note that Total Dominating Set is NP-complete on $K_{1,5}$-free split graphs [24] and polynomial time solvable on $K_{1,4}$-free split graphs [20]. Also, Total Dominating Set is NP-complete on $K_{1,3}$-free graphs [14] and an optimal total dominating set of a chordal bipartite graph can be found in polynomial time [3]. The study on open packing of graphs was initiated by Henning and Slater [9], and Rall [18] proved that for every non-trivial tree $T$, $\gamma_{t}(G)=\rho^{o}(G)$. In a recent work [21] (yet to be published), we extended this result by proving that the total domination number and the open packing number are equal when the underlying graph is a chordal bipartite graph with no isolated vertices. It is also known that Open Packing is NP-complete (NPC in short) on split graphs (equivalently, the class of $\{2K_{2},C_{4},C_{5}\}$-free graphs) [19] and planar bipartite graphs of maximum degree three (a subclass of $K_{3}$-free graphs) [21]. In this work, we complete the study of Open Packing on $H$-free graphs for every graph $H$ with at least three vertices by proving the following theorems.
@@
We also show that the bounds on $\rho^{o}(G)$ given in (ii) are tight. Moreover, we use the bounds on the total domination number in (a) and (b) to show that the results equivalent to those of Theorems 1 and 2 hold for Total Dominating Set in $H$-free graphs as well (see Theorem 9). This implies that Total Dominating Set and Open Packing are of the same nature in the view of classical complexity (P vs NPC) in $H$-free graphs for every graph $H$. Table 1 provides a comparative study of complexity (P vs NPC) between Total Dominating Set and Open Packing in certain classes of graphs. The rest of the results in this article are motivated by Table 1 and the question, ‘Does there exist a graph class where these two problems ‘differ’ from each other in the view of classical complexity (P vs NPC)?’
@@
We answered this question in the affirmative (see Table 2) with the following list of dichotomy results in subclasses of split graphs.
@@
    Open Packing is NP-complete on $K_{1,r}$-free split graphs for $r\geq 4$, and is polynomial time solvable for $r\leq 3$.
@@ 2 Preliminaries
We follow West [23] for terminology and notation. The graphs considered in this work are simple and undirected unless specified otherwise. Given a graph $G(V,E)$, let $n$ and $m$ denote the number of vertices and the number of edges in $G$, respectively. Given a vertex $x\in V(G)$, the (open) neighbourhood of $x$ in $G$ is defined as $N_{G}(x)=\{y\in V(G)\,:\,xy\in E(G)\}$, and let the degree of a vertex $x$ in $G$ be $deg_{G}(x)=|N_{G}(x)|$. The closed neighbourhood of a vertex $x$ in $G$ is defined as $N_{G}[x]=\{x\}\cup N_{G}(x)$. We use $N(x)$, $N[x]$ and $deg(x)$ instead of $N_{G}(x)$, $N_{G}[x]$ and $deg_{G}(v)$ respectively, when there is no ambiguity on $G$. A vertex $x$ in $G$ is called an isolated vertex in $G$, if $N_{G}(x)=\emptyset$. Given $U\subseteq V(G)$, the subgraph of $G$ induced by $U$ is denoted as $G[U]$. Given a graph $H$, $G$ is said to be $H$-free if no induced subgraph of $G$ is isomorphic to $H$. For a vertex $x\in V(G)$, let $E_{G}(x)$ denote the set of all edges incident on $x$, and for an edge $e\in E(G)$, let $V_{G}(e)$ denote the end vertices of $e$ in $G$. Note that for $u\in V(G)$ and $e\in E(G)$, the edge $e\in E_{G}(u)$ if and only if $u\in V_{G}(e)$. For a graph $G$, the line graph $L(G)$ of $G$ is a graph with $V(L(G))=E(G)$ and two elements $e,e^{\prime}\in V(L(G))$ are adjacent in $L(G)$ if and only if $V_{G}(e)\cap V_{G}(e^{\prime})\neq\emptyset$. Given two (not necessarily disjoint) graphs $H$ and $H^{\prime}$, the graph union $H\cup H^{\prime}$ is defined as $V(H\cup H^{\prime})=V(H)\cup V(H^{\prime})$ and $E(H\cup H^{\prime})=E(H)\cup E(H^{\prime})$. If $V(H)$ and $V(H^{\prime})$ are not disjoint, then we call the graph union $H\cup H^{\prime}$ as merger of the graphs $H$ and $H^{\prime}$. For $p\in\mathbb{N}\cup\{0\}$ and a graph $H$, the graph $pH$ is defined as the union of $p$ disjoint copies of $H$. Given two disjoint graphs $G$ and $H$, we say that (i) a vertex $x$ of the graph $G$ is replaced by the graph $H$ if a graph $G^{\prime}$ is constructed with $V(G^{\prime})=(V(G)\setminus\{x\})\cup V(H)$ and $E(G^{\prime})=(E(G)\setminus\{xy:y\in V(G)\text{ and }xy\in E(G)\})\cup E(H)\cup\{wy:w\in V(H),y\in(V(G)\setminus\{x\})\text{ and }xy\in E(G)\}$, and (ii) a vertex $x$ of the graph $G$ is identified with a vertex $y$ of the graph $H$, if the vertex $y$ in $H$ is relabelled as $x$ and the graphs $G$ and $H$ are merged.
@@
A vertex subset $C$ of $G$ is called a clique in $G$ if every pair of distinct vertices in $C$ is adjacent in $G$. A vertex subset $I$ of $G$ is called an independent set in $G$ if no pair of vertices in $I$ are adjacent in $G$. The size of a largest independent set in $G$ is called the independence number of $G$, and is denoted by $\alpha(G)$. Let $P_{n}$, $C_{n}$ and $K_{n}$ denote the path, cycle and complete graph on $n$ vertices, respectively. A set $D\subseteq V(G)$ is called a dominating set in $G$, if every vertex in $V(G)\setminus D$ is adjacent to some vertex in $D$. An edge subset $F$ of a graph $G$ is called a matching in $G$, if $V_{G}(e)\cap V_{G}(e^{\prime})=\emptyset$ for every pair of distinct $e,e^{\prime}\in F$. The cardinality of a largest matching in $G$ is called the matching number, $\alpha^{\prime}(G)$, of $G$.
@@ 3 $H$-free Graphs
We dedicate this section to prove the dichotomy result on Open Packing stated in Theorem 2. Observation 1 helps us prove the necessary part of Theorem 2.
@@
For $p\geq 4$, let $H$ be a graph on $p$ vertices such that $H\notin\{P_{4}\cup(p-4)K_{1},P_{3}\cup(p-3)K_{1},K_{2}\cup(p-2)K_{1},pK_{1}\}$. Then, $H$ contains at least one of $K_{3},2K_{2},C_{4},K_{1,3}$, or $C_{5}$ as an induced subgraph.
@@
Open Packing is NP-complete on (i) $2K_{2}$-free graphs, (ii) $C_{4}$-free graphs, (iii) $C_{5}$-free graphs, and (iv) triangle-free graphs because of the facts that Open Packing is NP-complete on (a) split graphs (i.e., the class of $\{2K_{2},C_{4},C_{5}\}$-free graphs) [19] and (b) bipartite graphs (a subclass of $K_{3}$-free graphs) [21, 22] and (c) Open Packing is in NP in every subclass of simple graphs.
Next, we prove the following results on $K_{1,3}$-free graphs: (i) Open Packing is NP-complete, (ii) Max-Open Packing is hard to approximate within a factor of $(N^{(\frac{1}{2}-\delta)})$ for any $\delta>0$ unless P = NP (where $N$ denotes the number of vertices in a $K_{1,3}$-free graph), and (iii) Open Packing parameterized by solution size is W[1]-complete.
@@ 3.1 $K_{1,3}$-free Graphs
Step 1: Replace each edge $e=uu^{\prime}$ in $G$ with a three vertex path $ueu^{\prime}$ in $G^{\prime}$. Step 2: For every $u\in V(G)$, make $E_{G}(u)$ a clique in $G^{\prime}$. Step 3: For every vertex $u_{i}\in V(G)$ with exactly one edge, say $e$ incident on it in $G$, introduce a vertex $v_{i}$, and two edges $u_{i}v_{i}$ and $v_{i}e$ in $G^{\prime}$.
@@
We can observe that in the graph $G^{\prime}$, for every vertex $\text{for every vertex }u_{i}\in V(G)$,
@@
The above equations imply that, by Steps 2 and 3 of Construction 1, $N_{G^{\prime}}(u_{i})$ and $N_{G^{\prime}}(v_{i})$ are cliques in $G^{\prime}$. Also, by Construction 1, for every $e\in E(G)$ with $V_{G}(e)=\{u,u^{\prime}\}$,
@@
This implies that $N_{G^{\prime}}(e)$ is a union of two cliques in $G^{\prime}$. We know that a graph $H$ is $K_{1,3}$-free if and only if $N_{H}(x)$ is a union of at most two cliques for all $x\in V(H)$. So, Equations (1), (2), and (3) imply that $G^{\prime}$ is $K_{1,3}$-free.
Next, we prove Claims 1-5, which are helpful in completing the proof of guarantee of Construction 1.
@@
Let $H$ be a graph, and let $U=\{x,y,z\}$ be a vertex subset of $H$ that induces a triangle in $H$. Then, for any open packing $S$ in $H$, $|S\cap U|\leq 1$.
@@
Let $u_{i}\in V(G)$ such that $deg_{G}(u_{i})=1$, and let $S$ be an open packing in $G^{\prime}$ with $v_{i}\in S$. Then, (i) $u_{i}\notin S$, (ii) $S^{\prime}=(S\setminus\{v_{i}\})\cup\{u_{i}\}$ is an open packing in $G^{\prime}$, and (iii) $|S|=|S^{\prime}|$.
Since $deg_{G}(u_{i})=1$, there exists an edge $e\in E(G)$ such that $E_{G}(u_{i})=\{e\}$. Then, $\{u_{i},v_{i},e\}$ induces a triangle in $G^{\prime}$ with $N_{G^{\prime}}(u_{i})=\{e,v_{i}\}$ and $N_{G^{\prime}}(v_{i})=\{e,u_{i}\}$. Therefore, by Observation 2, for any open packing $S_{1}$ in $G^{\prime}$, $|S_{1}\cap\{u_{i},v_{i},e\}|\leq 1$. Thus, $v_{i}\in S$ implies that $u_{i},e\notin S$. This proves (i) of Claim 2. Next, we prove (ii). Since $N_{G^{\prime}}(u_{i})=\{e,v_{i}\}$ and $N_{G^{\prime}}(v_{i})=\{e,u_{i}\}$, for any vertex $x\in S\setminus\{v_{i}\}$, $x$ and $v_{i}$ have no common neighbour in $G^{\prime}$, which implies that $x$ and $u_{i}$ have no common neighbour in $G^{\prime}$. This, together with the fact that $S\setminus\{v_{i}\}$ is an open packing in $G^{\prime}$, implies that $S^{\prime}=(S\setminus\{v_{i}\})\cup\{u_{i}\}$ is an open packing in $G^{\prime}$. Note that (iii) follows from (i).
@@
Let $S$ be an open packing in $G^{\prime}$ with $e=u_{\ell}u_{j}\in S\cap E(G)$ for some $1\leq\ell<j\leq n$. Then, (i) $u_{\ell},u_{j}\notin S$, (ii) $S_{\ell}=(S\setminus\{e\})\cup\{u_{\ell}\}$ and $S_{j}=(S\setminus\{e\})\cup\{u_{j}\}$ are open packings in $G^{\prime}$, and (iii) $|S|=|S_{\ell}|=|S_{j}|$.
Since $e\in S$, $u_{\ell},u_{j}\notin S$ because of Claim 3. This proves (i). Also, if $v_{\ell}\in V(G^{\prime})$, then $v_{\ell}\notin S$ because of Observation 2 and the fact that $\{v_{\ell},e,u_{\ell}\}$ induces a triangle in $G^{\prime}$. Similarly, if $v_{j}\in V(G^{\prime})$, then $v_{j}\notin S$. Next, we prove that $S_{\ell}=(S\setminus\{e\})\cup\{u_{\ell}\}$ is an open packing in $G^{\prime}$. Note that $S\setminus\{e\}$ is an open packing in $G^{\prime}$. So, if $S_{\ell}$ is not an open packing in $G$, then there exists a vertex $z\in S\setminus\{e\}$ such that $N_{G^{\prime}}(z)\cap N_{G^{\prime}}(u_{\ell})\neq\emptyset$. Let $y\in N_{G^{\prime}}(z)\cap N_{G^{\prime}}(u_{\ell})$. Note that $zu_{\ell}\notin E(G^{\prime})$, else $u_{\ell}\in N_{G^{\prime}}(z)\cap N_{G^{\prime}}(e)$, which would be a contradiction to $e,z\in S$ (an open packing in $G^{\prime}$). Similarly, $zu_{j}\notin E(G^{\prime})$. If $y=e$, then since $e=y\in N_{G^{\prime}}(z)$, we have $z\in(N_{G^{\prime}}(e)\setminus(N_{G^{\prime}}[u_{j}]\cup N_{G^{\prime}}[u_{\ell}]))$, a contradiction to Equation (3) (see Fig. 2). Therefore, $y\neq e$. This implies that $ye\in E(G^{\prime})$, since $y,e\in N_{G^{\prime}}(u_{\ell})$, which is a clique in $G^{\prime}$. Then, $y\in N_{G^{\prime}}(z)\cap N_{G^{\prime}}(e)$, a contradiction to $S$ being an open packing in $G^{\prime}$. Therefore, $S_{\ell}=(S\setminus\{e\})\cup\{u_{\ell}\}$ is an open packing in $G^{\prime}$. A similar argument holds for $S_{j}$. Also, note that (iii) follows from (i).